1-Let n X1, X2 ....Xn be random sample of size n from a distribution with probability

density function f(x,q)=qxq-1

0, else where.

Obtain a maximum likelyhood

estimator of q.



2- Let , X1, X2 Xn be independently and identically distributed b(1, p) random

variables. Obtain a confidence internal for p using Chebychev’s inequality


1
Expert's answer
2021-05-07T09:11:54-0400
  1. The given Probability density function is-

"f(x,q)=qx^{q-1}"


Taking log on both sides-

"log f(x,q)=logqx^{q-1}\\logf(x,q)=logq+(q-1)logx"


Log liklihood function is given by-

"L(X_1,X_2,X_3,...,X_n)=logf(x,q)=logq+(q-1)log x"


The maximum liklihood estimator of q-

"\phi_{ML}=\phi_{ML}(X_1,X_2,.....,X_n)"


2.Confidence interval by chebychev,s inequality is-


"P(|\dfrac{1-p}{n}|)<\dfrac{\sigma}{E^2}"


"P(|(X-E[X])|\ge k\sigma)\le \dfrac{1}{k^2}"